# AttributionLab: Faithfulness of Feature Attribution Under Controllable Environments

### v1

@@ 5 Faithfulness test of attribution methods in AttributionLab
In this section, we deploy the designed environment to test the faithfulness of attribution methods and analyze their various aspects. To provide an intuition about the behavior of various attribution methods, we visualize them in AttributionLab in Figure 4. This figure reveals a lack of consensus among attribution methods. To quantitatively test the alignment between attribution maps and ground truth masks, we use precision, recall, and $F_{1}$-score. However, since attribution maps usually have continuous pixel values, we adjust the metrics accordingly. Given a set of features $\mathcal{F}$ indexed by the index set $\mathcal{J}=\{j\in\mathbb{N}^{+}|1\leq j\leq|\mathcal{F}|\}$, let $a_{j}$ be the attribution value of the $j^{th}$ feature, and $g_{j}$ be the ground truth value of the $j^{th}$ feature, the (soft) precision is calculated as $Pr(\mathcal{F})=\sum_{j\in\mathcal{J}}|a_{j}\cdot g_{j}|/\sum_{j\in\mathcal{J}}|a_{j}|$, while the recall is defined as $Re(\mathcal{F})=\sum_{j\in\mathcal{J}}|a_{j}\cdot g_{j}|/\sum_{j\in\mathcal{J}}|g_{j}|$. We normalize the attribution maps to $[-1,1]$ to constrain the recall within the range of $[0,1]$, as the ground truth masks are binary. Given the precision and recall, we can easily calculate the $F_{1}$-score as well. In multi-class classification, signed ground truth information is available, outlining the positive and negative contributions of features to the target class. To test attribution methods with signed ground truth, we separately compute the precision, recall, and $F_{1}$-score of the positive and negative portions of the attribution maps using their corresponding ground truth masks. Furthermore, we test the entire attribution map using an unsigned union of both positive and negative ground truth: the overall ground truth. This test takes into consideration all features that contribute to decision-making, ignoring the sign of attribution values.
Furthermore, we employ these attribution methods on an ImageNet (Deng et al., 2009)-pretrained VGG16 (Simonyan & Zisserman, 2015) and check if the test result can generalize to the real world. Specifically, we present visual samples from various attribution methods to affirm their consistent behaviors as observed in synthetic setups.
Figure 4: Attributions in AttributionLab. The leftmost image visualizes the sampled data. The second leftmost image shows the ground truth attribution, where red and blue denote positive and negative attribution to the target class, respectively. The sample is randomly selected. More samples can be found in Appendix F
5.1 Faithfulness test of DeepSHAP
@@
Figure 5: (a) Faithfulness test of DeepSHAP in AttributionLab. (b) Visual example of DeepSHAP in synthetic and real-world environments. According to (a) and (b), DeepSHAP correctly highlights the foreground pixels. Nevertheless, it assigns both positive and negative attribution to these pixels, even when they have similar colors and close spatial locations. (c) Test result of LIME in AttributionLab. (d) Visual example of LIME in synthetic and real-world environments. Results in (c) and (d) show that the performance of LIME strongly depends on the segmentation result. [(Pos.) denotes the faithfulness test on positive attribution, while (Neg.) stands for the test on negative attribution, and Over. means the test on overall attribution neglecting the sign of attribution. Pre. and Rec. denote precision and recall, respectively.]
DeepSHAP (Lundberg & Lee, 2017) is designed to identify both positively and negatively contributing features. Hence, we test its faithfulness by comparing the positive attribution with positive ground truth features (GT), negative attribution with negative GT, and overall attribution with overall GT by only considering the magnitude of attribution values. Figure 4(a) shows the test result of DeepSHAP in the synthetic environment. The overall precision, recall, and $F_{1}$-score reveal that DeepSHAP performs well in locating the contributing features when we disregard the sign of attribution. However, the low precision and recall of positive and negative attribution suggest that DeepSHAP encounters difficulty in discerning whether a feature contributes positively or negatively to the target class. We further corroborated this issue by the visual results in Figure 4(b), which show the phenomenon in both synthetic and real-world scenarios. In the ImageNet example shown in the image, we observe both positive and negative attribution on the foreground object, especially on some pixels that are closely located and have similar colors. However, most of these highlighted pixels are likely to be relevant to the target class of the image (though we can’t ascertain which pixels the model is using). DeepSHAP is shown not to conform with Shapley value axioms (Sundararajan & Najmi, 2020). However, this claim does not tell us how faithful DeepSHAP is. Our results suggest that DeepSHAP can be used to identify all relevant features for the model without considering the sign of attribution values.
5.2 Faithfulness test of LIME
LIME (Ribeiro et al., 2016) requires the input image to be segmented into superpixels, it then treats all pixels within a superpixel as a single feature. Consequently, the resulting attribution map can be influenced by the segmentation step. To investigate the impact of segmentation, we utilize both Quickshift (Vedaldi & Soatto, 2008) and Felzenszwalb (Felzenszwalb & Huttenlocher, 2004) segmentation algorithms and test the faithfulness of the resulting attributions. Figure 4(c) reveals noticeable difference between the outcomes derived from these two segmentation techniques. Figure 4(d) provides additional visual evidence of these differences in real-world scenarios. In LIME, the attribution of a superpixel conveys the aggregate predictive influence of that segment. This “averaged” attribution is then broadcasted across all pixels within the superpixel. A more fine-grained attribution map requires a finer segmentation process. Hence, LIME requires prior knowledge regarding the ground truth features, that is, which pixels belong together and which ones are independent.
5.3 Faithfulness test of Integrated Gradients (IG), and Occlusion
@@
Previous research (Sturmfels et al., 2020) has revealed that IG (Sundararajan et al., 2017) is sensitive to the choice of baseline. While testing IG with a proper baseline remains a challenge in the real world, our controlled experimental setup provides the unique advantage of access to the true baseline. We introduce this baseline-accurate variant of IG as $\text{IG}^{*}$, which uses the true baseline corresponding to background color $(20,20,20)$ in synthetic images. Despite that the visual difference between the true baseline and the widely used baseline $(0,0,0)$ is small, Figure 5(a) reveals a significant enhancement in the precision of $\text{IG}^{*}$. We see that sensitivity to baseline is not an ignorable issue. Analogous to $\text{IG}^{*}$, we introduce $\text{Occlusion}^{*}$ that employs the true baseline. Figure 5(b) also demonstrates a notably improved precision of $\text{Occlusion}^{*}$. Figure 5(c) and Figure 5(d) further illustrate the sensitivity of these methods to the choice of baseline. Based on these empirical findings, we underscore that one potential direction for enhancing these methods is the determination of an accurate baseline, even finding the baseline for each image separately (e.g., through generative models). Additionally, the performance of $\text{Occlusion}^{*}$ still remains less than optimal. One reason is that Occlusion’s efficacy is influenced by the size of the sliding window used for pixel perturbation. Determining an appropriate perturbation area requires prior knowledge of both the dataset and the model structure, which is sometimes challenging to acquire in real-world applications.
5.4 Faithfulness test of GradCAM, GuidedBP, ExPerturb, IBA
(a) Faithfulness test results
(b) Visual examples in synthetic setting and ImageNet
Figure 7: Faithfulness test and visual examples for GradCAM, GuidedBP, ExPerturb, and IBA. Some effectively identify positively contributing features, but none successfully discern negatively contributing features, inherent to their design limitations. GradCAM and IBA exhibit blurriness upon resizing, and the performance of ExPerturb is markedly influenced by the perturbation area.
GradCAM (Selvaraju et al., 2017) and GuidedBP (Springenberg et al., 2015) demonstrate high recalls on identifying positively contributing pixels. However, they do not identify pixels that negatively contribute to the target class, as evidenced by Figure 4 and Figure 7. This is because these methods typically initiate backpropagation from the logits before the softmax. The negatively contributing pixels do not influence backpropagation from the target class logit. Additionally, GradCAM resizes the attribution map to match the dimensions of the input. This resizing operation introduces blurriness (Figure 6(b)), which consequently diminishes its precision. As depicted in Figure 6(a), ExPerturb (Fong et al., 2019) displays higher recall compared to its precision. This reduced precision can be attributed to the improper choice of the perturbation area constraint. In addition, Figure 4 and Figure 6(b) show that IBA (Schulz et al., 2020) can locate predictive features. However, the sigmoid used in the information bottleneck leads to only non-negative attribution.
6 Faithfulness of evaluation metrics
There is another perspective through which faithfulness can be evaluated. We can perturb pixels in the image based on attribution values and observe how the output of the neural network changes. If the output change is aligned with the attribution values, then the attribution is considered faithful from this perspective. For instance, Insertion and Deletion (Samek et al., 2016) progressively insert or remove pixels, starting from the pixel with the highest attribution to the lowest attribution. In contrast to perturbing pixels according to a specific order, Sensitivity-N (Ancona et al., 2018) randomly selects $N$ pixels to perturb, after that measuring the correlation between the change in prediction and the sum of attribution of the perturbed pixels.
However, there is a fundamental issue lurking within this perspective on faithfulness. A significant output change can result from the network not having seen the new input resulting from perturbation rather than the perturbed pixel being important. In the controlled environment of AttributionLab, we can observe how sensitive these evaluation metrics are to Unseen Data Effect. That is when the pixels are perturbed during evaluation to another value for which the network is not programmed, a scenario that happens easily in practice. We conduct experiments using these metrics on two AttributionLab setups to assess how this phenomenon (Unseen Data Effect) affects their performance. Figure 8 shows that the order of feature attribution methods in terms of performance changes between the two scenarios for both insertion and Sensitivity-N metrics. To confirm this observation, we compare the performance rankings of attribution methods in these metrics with performance rankings established from the ground truth. Therefore, we first rank the attribution methods using the $F_{1}$-score derived from positive ground truth attribution. Subsequently, we compute Spearman’s rank correlation between the rankings, as demonstrated in Table 1. Additional experimental results are shown in Appendix F. The first row of Table 1 reveals that, in the presence of Unseen Data Effect, the metrics display substantially less consistency with the ground-truth-based evaluation. This inconsistency arises because Unseen Data Effect can result in unexpected predictions and inaccurate estimation of changes in model output. Hence, using model output scores as metrics may not accurately report the performance of attribution methods in the presence of Unseen Data Effect.
@@ 6 Faithfulness of evaluation metrics
Figure 8: Sensitivity of perturbation-based evaluation metrics to Unseen Data Effect. We test these metrics on the two AttributionLab setups to observe the effect of this phenomenon on these metrics. (a) and (b) show the insertion metric results with and without Unseen Data Effect, respectively. (c) and (d) show the sensitivity-N metric results with and without the Unseen Data Effect. (in all metrics, a higher curve is better.) It is observed that with the presence of the Unseen Data Effect, the order of attribution methods changes in these metrics.
@@
Table 1: Spearman’s rank correlation. The rankings of attribution methods are measured on perturbation-based metrics and the designed ground truth, respectively. In the presence of the Unseen Data Effect, these metrics show significant deviation from ground-truth-based evaluations.
Model has ”Unseen Data Effect” Insertion Deletion Sensitivity-N
@@ 7 Conclusion and future work
In this work, we propose a novel, controlled laboratory setup for testing feature attribution explanations. The crux of our approach is the implementation of a paired design, i.e., manually programming the neural network and designing the dataset. We test feature attribution explanations by simulating various conditions such as inappropriate baseline values for attribution methods and different segmentation masks as input of attribution methods, demonstrating their significant impact on attribution performance. We observe that the sensitivity to the baseline of Shapley value methods (SHAP and Integrated gradients) due to Unseen Data Effect is not an ignorable issue. We show that the Unseen Data Effect problem inherent in perturbation-based evaluations can have negative effects on evaluation findings. Our proposed synthetic environment can empower future research by identifying potential failure modes of attribution methods in a trustable, controlled environment. This could help researchers address these issues before deploying these methods in real-world settings, thus contributing to the development of more reliable and effective attribution methods. There are several potentials for future exploration. The community could develop additional settings to address a broader spectrum of potential issues in attribution methods. In addition, constructing more complex synthetic models and datasets would better approximate real-world scenarios. Furthermore, expanding this methodological approach to other data modalities and architectural designs could provide a more comprehensive understanding of feature attribution explanation methods.

### v2

@@ 4 Faithfulness test of attribution methods
In this section, we deploy the designed environment to test the $(\gamma,\mathcal{F})$-model agnostic faithfulness of attribution methods and analyze their various aspects. We first visualize the attribution methods in AttributionLab in Figure 2. This figure reveals a lack of consensus among attribution methods, even in our synthetic environments. To quantitatively test the alignment between attribution maps and ground truth masks, we use precision, recall, and $F_{1}$-score. However, since attribution maps usually have continuous pixel values, we adjust the metrics accordingly. Given a set of features $\mathcal{S}$ indexed by the index set $\mathcal{J}=\{j\in\mathbb{N}^{+}|1\leq j\leq|\mathcal{S}|\}$, let $a_{j}$ be the attribution value of the $j^{th}$ feature, and $g_{j}$ be the ground truth value of the $j^{th}$ feature, the (soft) precision is calculated as $Pr(\mathcal{S})=\sum_{j\in\mathcal{J}}|a_{j}\cdot g_{j}|/\sum_{j\in\mathcal{J}}|a_{j}|$, while the recall is defined as $Re(\mathcal{S})=\sum_{j\in\mathcal{J}}|a_{j}\cdot g_{j}|/\sum_{j\in\mathcal{J}}|g_{j}|$. We normalize the attribution maps to $[-1,1]$ to constrain the recall within the range of $[0,1]$, as the ground truth masks are binary. Given the precision and recall, we can easily calculate the $F_{1}$-score as well. In multi-class classification, signed ground truth information is available, outlining the positive and negative contributions of features to the target class. To test attribution methods with signed ground truth, we compute the precision, recall, and $F_{1}$ by comparing the positive attribution with positive ground truth features (GT), negative attribution with negative GT, respectively. Furthermore, we test the entire attribution map using an unsigned union of both positive and negative ground truth: the overall ground truth. This test takes into consideration all features that contribute to decision-making, ignoring their direction of contribution. To measure $(\gamma,\mathcal{F})$-model agnostic faithfulness in Definition 2.1, we employ the $F_{1}$-score as the faithfulness measurement. We use $\gamma=0.5$ and consider $(0.5,\mathcal{F})$-model agnostic faithfulness as the criterion for our faithfulness test in the main text.
(a) Faithfulness test
(b) Visual examples
@@
Figure 5: Faithfulness test and visual examples of DeepSHAP. (a) Faithfulness test of DeepSHAP. (b) Visual examples in synthetic and real-world environments. According to (a) and (b), DeepSHAP correctly highlights the foreground pixels. However, it assigns both positive and negative attribution to these pixels, even when they have similar colors and close spatial locations.
(a) Faithfulness test
(b) Visual examples
Figure 6: LIME with diverse segmentation priors. (a) Faithfulness test of LIME in AttributionLab. (b) Visual examples of LIME in synthetic and real-world environments. Results (a,b) show that the faithfulness of LIME depends on the segmentation prior.
Additionally, we employ these attribution methods on an ImageNet (Deng et al., 2009)-pretrained VGG16 (Simonyan & Zisserman, 2015) to validate our findings from AttributionLab in real-world trained models. Specifically, we present visual samples from various attribution methods to affirm their consistent behaviors as observed in synthetic setups.
4.1 DeepSHAP faithfulness analysis
@@
— When considering all relevant features (discarding their sign), DeepSHAP passes the faithfulness test.
DeepSHAP (Lundberg & Lee, 2017) is designed to identify both positively and negatively contributing features. Figure 4(a) shows the test result of DeepSHAP in the synthetic environment. The overall precision, recall, and $F_{1}$-score reveal that DeepSHAP effectively locates the contributing features when we disregard the sign of attribution. Subsequently, DeepSHAP passes the faithfulness test with $\gamma=0.5$. However, the low precision and recall of positive and negative attribution suggest that DeepSHAP encounters difficulty in discerning whether a feature contributes positively or negatively to the target class. When we consider positive or negative attribution only, DeepSHAP fails the faithfulness test in both cases under the same $\gamma$. We further corroborated this issue in Figure 4(b), which shows the phenomenon in both synthetic and real-world scenarios. In the ImageNet example shown in the image, we observe both positive and negative attribution on the foreground object, especially on some pixels that are closely located and have similar colors. However, most of these highlighted pixels are likely to be relevant to the decision-making (though we cannot ascertain which pixels the model actually uses). Our results suggest that DeepSHAP can be used to identify relevant features for the model, but is not effective in determining the direction of the contribution, although it is designed to do so.
4.2 LIME faithfulness analysis
— Using the right segmentation prior, LIME passes the faithfulness test.
LIME (Ribeiro et al., 2016) requires the input image to be segmented into superpixels, it then treats all pixels within a superpixel as a single feature. Consequently, the resulting attribution map can be influenced by the segmentation step. To investigate the impact of segmentation, we utilize Quickshift (Vedaldi & Soatto, 2008), Felzenszwalb (Felzenszwalb & Huttenlocher, 2004), and SAM (Kirillov et al., 2023) segmentation algorithms and test the faithfulness of the resulting attributions. Figure 5(a) reveals a noticeable difference between the outcomes derived from these segmentation techniques. Furthermore, LIME using Felzenszwalb fails to pass the faithfulness test with $\gamma=0.5$. Lime using Quickshift fails the faithfulness test with $\gamma=0.6$. However, LIME with SAM can pass the test for both values of $\gamma$. Figure 5(b) provides additional visual evidence of these differences in real-world scenarios. In LIME, the attribution of a superpixel conveys the aggregate predictive influence of that segment. This “averaged” attribution is then broadcast across all pixels within the superpixel. A more fine-grained attribution map requires a finer segmentation process. Therefore, a promising way to improve LIME is to find a better segmentation prior, that is, which pixels belong together and which ones are independent.
4.3 IG and Occlusion faithfulness analysis
— IG and Occlusion pass faithfulness tests only with a proper baseline.
Previous research (Sturmfels et al., 2020) has revealed that IG (Sundararajan et al., 2017) is sensitive to the choice of baseline. While testing IG with a proper baseline remains a challenge in the real world, our controlled experimental setup provides the unique advantage of access to the true baseline. To study the effect of the baseline, we apply two settings. Here, $\text{IG}^{*}$ uses the predefined true baseline corresponding to background color $(20,20,20)$ in synthetic images, while IG employs a slightly incorrect baseline $(0,0,0)$ as default in many attribution applications. Figure 6(a) reveals a significant enhancement in the precision of $\text{IG}^{*}$ in our synthetic setting. Specifically, IG fails the faithfulness test with $\gamma=0.5$, while $\text{IG}^{*}$ passes. Analogous to $\text{IG}^{*}$, we introduce $\text{Occlusion}^{*}$ that employs the true baseline. Figure 6(b) also demonstrates a notably improved precision of $\text{Occlusion}^{*}$. Figure 6(c) and Figure 6(d) further illustrate the sensitivity of these methods to the choice of baseline. Faithfulness test results show that IG and Occlusion are sensitive to the choice of baselines. In addition, they fail the faithfulness test given incorrect baselines. As methods like IG have adequate theoretical support, our findings underscore the importance of correct baselines for these methods. Future research on these methods should focus on finding correct baselines to enhance their performance.
(a) Faithfulness test
(b) AttributionLab and ImageNet samples
Figure 8: Faithfulness test and visual examples for GradCAM, ExPerturb, and IBA. All methods show promising recall in picking up positive features. ExPerturb has low precision (optimized mask can include other features). IBA is relatively the most faithful in identifying positive relevant features.
4.4 GradCAM, IBA and ExPerturb
— IBA and GradCAM pass the faithfulness test in identifying positive relevant features.
These methods have different mechanisms: (GradCAM uses activations, ExPerturb uses mask optimization on input, and IBA masks the latent activations). However, they have two similar properties. First, they are not granular and are smooth (approximate) by design. Second, they are explicitly designed to provide positive contributions only. Thus, they are not capable of reflecting negatively relevant features. Therefore, we only analyze faithfulness to positively relevant features (Figure 8), and to be fair to their granularity issue, we use smooth ground truth. GradCAM (Selvaraju et al., 2017) shows high recalls on identifying positively contributing pixels. It resizes the attribution map to match the dimensions of the input. This resizing operation introduces blurriness (Figure 7(b)), which consequently diminishes its precision. As for ExPerturb (Fong et al., 2019), the optimized mask also involves features that are not positively contributing to the output. Thus, the precision is low, though the recall for finding positive features is high. IBA (Schulz et al., 2020) achieves high precision and recall similar to GradCAM. In conclusion, all three methods pass the faithfulness test for $(0.5,\mathcal{F})$-model agnostic faithfulness when we only consider positive contributions.
(a) W/ Unseen Data Effect
(b) W/o Unseen Data Effect
@@ 4.4 GradCAM, IBA and ExPerturb
Figure 9: Sensitivity of perturbation-based evaluation metrics to Unseen Data Effect. We test Insertion on the two AttributionLab setups. In (b) several lines are overlapped. With the presence of the Unseen Data Effect, the evaluation result changes. Detailed results of Deletion and Sensitivity-N are in Appendix G.
5 Sanity check of faithfulness evaluations
— We simulate the Unseen Data Effect in AttributionLab and observe that perturbation-based faithfulness evaluations are not reliable.
There is another perspective through which faithfulness can be evaluated. One can perturb pixels in the image based on attribution values and observe how the output of the neural network changes. If the output change is aligned with the attribution values, then the attribution is considered faithful from this perspective. For instance, Insertion and Deletion (Samek et al., 2016) progressively insert or remove pixels (Sensitivity-N (Ancona et al., 2018) is in Appendix B and G.5).
However, there is a fundamental issue lurking within this perspective on faithfulness. A significant output change can result from the network not having seen the new input resulting from perturbation, rather than the perturbed pixel being important. In the controlled environment of AttributionLab, we can observe how sensitive these evaluation metrics are to the Unseen Data Effect. That is when the pixels are perturbed during evaluation to another value for which the network is not trained, a scenario that happens easily in practice. We conduct experiments using these metrics on two AttributionLab setups to assess how this phenomenon (Unseen Data Effect) affects their performance. Figure 9 shows the order of feature attribution methods in terms of performance changes between the two scenarios for Insertion metrics. To confirm this observation, we compare the performance rankings ($F_{1}$-score derived from positive ground truth attribution) of attribution methods in these metrics with performance rankings established from the ground truth. Subsequently, we compute Spearman’s rank correlation between the rankings, as demonstrated in Table 2. Additional experimental results are shown in Appendix G. The first row of Table 2 reveals that, in the presence of Unseen Data Effect, the metrics display substantially less consistency with the ground-truth-based evaluation. This inconsistency arises because Unseen Data Effect can result in unexpected predictions and inaccurate estimation of changes in model output. Hence, using model output scores as metrics may not accurately report the performance of attribution methods in the presence of Unseen Data Effect.
@@ 5 Sanity check of faithfulness evaluations
Table 2: Spearman’s rank correlation. The attribution method rankings are measured on perturbation-based metrics and the designed ground truth. In the presence of the Unseen Data Effect, these metrics show significant deviation from the ground truth.
Model has Insertion Deletion Sensitivity-N
Unseen Data Effect
@@
6 Related work
Since the advent of deep neural networks (Krizhevsky et al., 2012), understanding how these complex models make predictions came under the spotlight (Simonyan et al., 2013). One approach is feature attribution. Initial efforts to perform feature attribution focused on linear models (Simonyan et al., 2013; Bach et al., 2015) and backpropagation (Zeiler & Fergus, 2014; Springenberg et al., 2015). Subsequently, more principled approaches emerged, such as backpropagating output differences with respect to a reference input (Bach et al., 2015; Shrikumar et al., 2017) and axiomatic methods (Sundararajan et al., 2017; Lundberg & Lee, 2017) inspired by the Shapley value (Shapley, 1953). Meanwhile, intuitive approaches probing internal states (Selvaraju et al., 2017; Chattopadhay et al., 2018; Schulz et al., 2020) and optimizing input masks were introduced (Fong et al., 2019; Zhang et al., 2021). With these advancements, feature attribution evaluation became a central question. Literature has proposed sanity checks testing whether attribution changes upon randomizing the network’s weights (Adebayo et al., 2018) or if it is used on a different output class (Sixt et al., 2020). It also evaluates faithfulness by analyzing networks’s output when perturbing input features based on their relevance (Samek et al., 2016; Ancona et al., 2018; Montavon et al., 2018; Hooker et al., 2019; Rong et al., 2022), or show theoretically if they are aligned with axioms (Sundararajan et al., 2017; Lundberg & Lee, 2017). Each reveals different aspects (Appendix B). A recent trend in evaluating faithfulness is designing datasets with known input-feature-output associations, enabling comparison between attribution and ground truth. Arras et al. (2022); Zhou et al. (2022); Khakzar et al. (2022); Agarwal et al. (2023). However, these works do not guarantee that the network uses the intended ground truth associations.
7 Limitation
We introduce an attribution faithfulness test by creating a synthetic environment, including a synthetic neural network and a synthetic dataset. Using our analysis in the controlled environment, we provide systematic suggestions for improving attribution methods. While this approach offers an effective means for sanity check and troubleshooting of attribution methods, it has several limitations. Despite our efforts to design a neural network with conventional architectures and scales, our model does not fully resemble a trained neural network. Hence, designing other models that further imitate trained model behaviors can expand the test scenarios. The current study is limited to vision tasks, Convolutional Neural Networks (CNNs), and MLPs. Studies of transformers and other data modalities are not discussed in this work. With the addition of a self-attention module, transformers can be realized in follow-up research.
@@ 8 Conclusion and future work
This work proposes a novel, controlled laboratory setup for testing feature attribution explanations. The crux of our approach is the implementation of a paired design, i.e., manually programming the neural network and designing the dataset. We leverage our synthetic environment for a novel faithfulness test that efficiently and accurately checks whether an attribution method is not model agnostically faithful. We test feature attribution explanations by simulating various conditions such as inappropriate baseline values for attribution methods and different segmentation masks as input of attribution methods, demonstrating their significant impact on attribution faithfulness. Our findings highlight the importance of correct settings of the existing attribution methods and encourage more exploration in finding adequate configurations than proposing new attribution methods. We show that the Unseen Data Effect problem inherent in perturbation-based evaluations can negatively affect evaluation findings. Our proposed synthetic environment can empower future research by identifying not faithful methods and studying potential failure modes of attribution methods in a trustable, controlled environment. This could help researchers address issues before deploying methods in real-world settings, thus contributing to the development of more reliable and effective attribution methods.
Impact statements
In this work, we introduce an evaluation framework designed specifically for feature attribution methods. There is a growing body of research focused on enhancing the explainability of large models in response to concerns surrounding their transparency. However, having wrong but positively affirming explanations, can have a large negative societal impact. Nonetheless, providing incorrect yet positively affirming explanations can potentially yield significant adverse societal consequences. A notable gap in the research community is the absence of reliable evaluation metrics, leading to various complications. For example, numerous feature attribution methods have been proposed that, in some cases, even yield inconsistent attribution results given identical inputs. Our approach guarantees its validity for evaluation purposes because of its fully synthetic nature. Our synthetic settings may appear primitive compared to real-world datasets and neural networks trained on such datasets. However, they provide a controlled laboratory environment, enabling a thorough examination of feature attribution methods prior to deployment. This facilitates a robust evaluation and refinement process for attribution methods. In this controlled environment we can identify the factors that render each attribution unfaithful, and provide a roadmap for large-scale analysis of these attributions.